### v2

# Spectra and pseudospectra in the evaluation of material stability

Abstract
We consider the dynamics of bodies with “active” microstructure described by vector-valued phase fields. For waves with time-varying amplitude, the associated evolution equation involves a matrix that can be non-normal, depending on the constitutive choices adopted for the microstructural actions associated with the phase field. In the presence of non-normality, the spectral analysis of such a matrix when (unknown or uncertain or time-dependent) constitutive parameters vary is generically not decisive to evaluate the material stability. We thus need to look at the pseudospectrum, that is, the set of all possible eigenvalues of matrices in an $\varepsilon$-neighborhood of the matrix of interest. As a case study, we look at quasicrystals and evaluate through spectra and pseudospectra the influence of a microstructural self-action on the material stability. This approach allows us to reliably determine thresholds where stable-unstable transitions in the material behavior occur.
keywords:
Microstructures; Phase fields; Waves; Quasicrystals; Pseudospectra; Material Stability; Continuum Mechanics
1 Introduction
In the dynamics of bodies with “active” microstructure described by vector-valued phase fields we consider waves with time-varying amplitude. Their behavior is governed by the interactions between standard and microstructural actions, the latter ones associated with the (independent) mechanisms described by the phase field. Such microstructural interactions are a microstress (first-neighbor contact interaction) and a bulk self-action; their balance supplements the balance of macroscopic momentum and is coupled with it. No peculiar microstructural inertia [5] is attributed to phenomena described by the phase field; only possible microstructural (diffusion-type) dissipation is considered.
In this multi-field setting, constitutive structures may involve—and often they actually do—unknown or uncertain parameters. Parametric analyses may thus reveal admissible ranges for material stability and thresholds towards the unstable behavior. Results may thus address experiments aiming at the quantification of these unknown parameters. In developing such analyses in the mechanics of complex bodies (those with “active” microstructure) the occurrence of non-normal matrices is not infrequent, as in the case treated here of time-dependent amplitude waves. With the adjective “non-normal” we indicate those matrices $A\in\mathbb{C}^{n\times n}$ such that their product with the complex conjugate transpose matrix does not commute, namely $AA^{\ast}\neq A^{\ast}A$. For such matrices, the spectral analysis alone is in general not decisive for checking material stability or the transition to unstable behavior. Rather, we need to call upon the notion of pseudospectrum as introduced first by Trefethen and Trummer with reference to Chebyshev and Legendre spectral differentiation matrices, adopted in performing the explicit time (numerical) integration of an hyperbolic initial boundary value problem [37], [36] (see also [14]). The basic idea is to examine the spectra of matrices lying within a ball or radius $\varepsilon$ around the non-normal matrix $A$. The resulting sets of eigenvalues constitute the so-called pseudospectrum and provide more accurate information on possible stable-unstable transition in the mechanical behavior [36].
As a concrete example, we perform explicit numerical analyses on wave propagation in quasicrystals. They are materials with perfect long-range order, but with no three-dimensional translation periodicity [32]. Their structure is made of atomic clusters distributed with a symmetry that is incompatible with a periodic tiling of the ambient space, for example icosahedral symmetry in $3D$-space [32], [31]. The quasi-periodic lattice is characterized by additional low-spatial-scale degrees of freedom, so-called phasons, here indicated by $\nu$ at $x$ and $t$. They are justified on geometrical basis. Indeed, if we expand in Fourier series the mass distribution of a $3D$ quasi-periodic lattice, the resulting wave vectors are six-dimensional [31]. More generally, a quasi-periodic lattice in a $n$-dimensional point space can be seen as a projection of a periodic lattice in a higher dimensional ambient onto a $nD$ incommensurate subspace. To visualize the circumstance, consider a $2D$ lattice with square symmetry; take in the plane a straight line inclined by an irrational angle with the lattice symmetry axes and project over it, orthogonally, the lattice atoms in a strip around the line (whatever be its thickness); the result is a quasi-periodic $1D$ lattice over the line, which is in this case the incommensurate subspace. A displacement in the higher-dimensional periodic lattice has thus a component in the incommensurate subspace and one orthogonal to it; the first projection of the displacement field is what is commonly called a phonon field, because it is associated with acoustic waves; the orthogonal displacement component to the incommensurate subspace leads to the field of phasons. This geometrical aspect suggested first a convenient description of the quasicrystal behavior under external actions as a higher-dimensional replica of the mechanics of simple bodies [12], [17]. However, if we consider the degrees of freedom represented by $\nu$ as those exploited by microscopic rearrangements that assure quasi-periodicity of the atomic lattice under given (possibly time-varying) boundary conditions, at continuum scale they are inner to every material element associated with a geometric point; so, $\nu$ itself has to be considered insensitive to relative rigid translations of observers in the physical space while, being a vector, it is sensitive to observer rotations. The circumstance detaches description of quasicrystal continuum mechanics from a higher dimensional replica of the standard format. Rather it falls within the general model-building framework for the mechanics of complex bodies (as shown in [22], [25]), in particular in the special case treated here, in which the phase field describing microstructural effects takes values in $\mathbb{R}^{3}$. In this general setting, balance equations are those of forces and those of microstructural interactions associated with the phase field; a balance of couples put together in the same expression standard and microstructural stresses, and a self-action at microstructure level. Although these balance equations can be postulated as first principles [4], they can be derived from an invariance principle together with the representation of contact actions in terms of standard and microstructural stresses. The principle mentioned can be the requirement of objectivity imposed to the external power alone [21] or the covariance (that is, structure-invariance under diffeomorphism-based observer changes) of the Clausius-Duhem inequality, written in terms of external power [24]. Resorting to first principles allows an appropriate control on the results unlike a mere use of analogies that can be surely useful but, in essence, are only hopes, although at times concrete ones.
As regards quasicrystals, the invariance techniques mentioned (and here we adopt the $SO(3)$ invariance of the external power alone) imply the possible emergence of a phason self-action $\mathbf{z}$ which can have both conservative (energetic-type) and dissipative components [25]. In treatments of quasicrystal elasticity based on analogy to simple bodies in a higher dimensional spaces [12] a conservative component of $\mathbf{z}$ is completely discarded with the obvious advantage of reducing the analysis to a replica of standard elasticity with the proviso of considering a higher-dimensional space. When phason friction is considered, a dissipative component of $\mathbf{z}$ comes into play [30]. The conservative component of $\mathbf{z}$ should play, however, a role in material stability. To explore the circumstance, we analyze the time-dependent amplitude wave propagation in different cases:
  1. 1)
    absence of a self-action, namely $\mathbf{z}=0$;
  2. 2)
    purely conservative character of the self-action, meaning that $\mathbf{z}$ equals the derivative of a free energy density $\psi$ with respect to $\nu$; a special choice is $\mathbf{z}=k_{0}\nu$, $k_{0}\in\mathbb{R}^{+}$;
  3. 3)
    purely dissipative self-action, so that $\mathbf{z}$ represents microstructural friction alone; thus, it is such that $\mathbf{z}\cdot\dot{\nu}\geq 0$, for any $\dot{\nu}$, an inequality that is compatible with the choice $\mathbf{z}=c\dot{\nu}$, $c\in\mathbb{R}^{+}$;
  4. 4)
    simultaneous presence of both conservative ($\mathbf{z}^{e}$) and dissipative ($\mathbf{z}^{d}$) components of the self-action, namely $\mathbf{z}=\mathbf{z}^{e}+\mathbf{z}^{d}$, where $\mathbf{z}^{e}=k_{0}\nu$ and $\mathbf{z}^{d}=c\dot{\nu}$.
By spectral and pseudospectral analyses of the system that governs the wave amplitude evolution, we determine for the experimentally unknown constitutive parameters possible thresholds for stable-unstable transition in the material behavior.
We refer all the following analyses to orthonormal frames of reference for the sake of simplicity. So, we do not distinguish between covariant and contravariant components of tensors; similarly we avoid distinguishing between transpose (indicated by a superscript $\top$) and formal adjoint (commonly denoted by an asterisk alone) of linear operators.
An interposed dot will eventually indicate the scalar product between elements of a linear space, whatever it be.
2 Basic fields
Consider two isomorphic copies of the $3D$ real space, namely ^$\hat{\mathbb{R}}^{3}$ and $\mathbb{R}^{3}$; the isomorphism $\hat{\iota}:\hat{\mathbb{R}}^{3}\longrightarrow\mathbb{R}^{3}$ is merely the identification map. In ^$\hat{\mathbb{R}}^{3}$ we select a fit region $\mathcal{B}$, a bounded simply connected domain endowed with a surface-like boundary oriented by the outward unit normal $\mathbf{n}$ to within a finite number of corners and edges. We adopt $\mathcal{B}$ as a macroscopic reference configuration. From it, deformations are orientation-preserving differentiable one-to-one maps $x\longmapsto y:=\tilde{y}(x)\in\mathbb{R}^{3}$. We indicate by $F$ the deformation gradient. To account for motions in a time interval $[0,\bar{t}\,]$, we extend $\tilde{y}$ to the space-time tube ×$\mathcal{B}\times[0,\bar{t}\,]$ so that we have $(x,t)\longmapsto y:=\tilde{y}(x,t)\in\mathbb{R}^{3}$ and we take $\tilde{y}(x,\cdot)$ to be differentiable also with respect to time at every $x\in\mathcal{B}$; a superposed dot will indicate the pertinent derivative. We thus define a displacement $\tilde{u}$ in the physical space as
$(x,t)\longmapsto u:=\tilde{u}(x,t)=\tilde{y}(x,t)-\hat{\iota}(x)\;.$
Thus, we have $F=I+\nabla u$, where $I$ is the shifter from ^$\hat{\mathbb{R}}^{3}$ to $\mathbb{R}^{3}$, with components $\delta^{i}_{\;A}$, where capital indices refer to coordinates in ^$\hat{\mathbb{R}}^{3}$, while the others to their counterparts in $\mathbb{R}^{3}$.
The velocity in Lagrangian representation is given by $\dot{y}=\frac{\mathrm{d}\tilde{y}(x,t)}{\mathrm{d}t}=\dot{u}$.
A differentiable phase field
$(x,t)\longmapsto\nu:=\tilde{\nu}(x,t)\in\tilde{\mathbb{R}}^{3}$
brings at macroscopic scale information on microstructural events at spatial scale $\lambda$. ~$\tilde{\mathbb{R}}^{3}$ is isomorphic to $\mathbb{R}^{3}$ but distinguished from it; the isomorphism is once again the identification. We indicate by $N$ its gradient, namely $N:=\nabla\nu$. The time rate of $\nu$ in Lagrangian representation is given by $\dot{\nu}=\frac{\mathrm{d}\tilde{\nu}(x,t)}{\mathrm{d}t}$.
The small strain regime is commonly defined by the condition $|\nabla u|\ll 1$. In such a regime we do not distinguish between ^$\hat{\mathbb{R}}^{3}$ and $\mathbb{R}^{3}$.
3 Invariance and balance
3.1 Observers and their changes
The prescription of reference frames over all the spaces used for representing the morphology of a body and its motion is what we consider to be an observer. Here such spaces are the reference one (namely ^$\hat{\mathbb{R}}^{3}$), the physical space $\mathbb{R}^{3}$, ~$\tilde{\mathbb{R}}^{3}$ where $\tilde{\nu}$ takes values, and the time interval. We adopt observer changes that leave invariant the time scale and $\hat{\mathbb{R}^{3}}$, while they are isometries in $\mathbb{R}^{3}$ and ~$\tilde{\mathbb{R}}^{3}$.
More precisely, if $y$ is a place evaluated by a first observer $\mathcal{O}$, another observer $\mathcal{O}^{\prime}$ records a counterpart of $y$ given by $y^{\prime}=a(t)+Q(t)(y-y_{0})+y_{0}$, where $t\longmapsto a(t)\in\mathbb{R}^{3}$ is a vector-valued smooth map depending only on time; $y_{0}$ is an arbitrary fixed point, and $t\longmapsto Q(t)\in SO(3)$ is an orthogonal tensor-valued smooth map depending only on time.
$\mathcal{O}$ records a velocity $\dot{y}$, which is $\dot{y}^{\prime}=\dot{a}+\dot{Q}(y-y_{0})+Q\dot{y}$ for $\mathcal{O}^{\prime}$. By pulling back $\dot{y}^{\prime}$ to $\mathcal{O}$, we get $\dot{y}^{\diamond}:=Q^{\mathsf{T}}\dot{y}^{\prime}$ given by
$\dot{y}^{\diamond}=\mathfrak{c}+q\times(y-y_{0})+\dot{y}\;.$
The vector $\mathfrak{c}:=Q^{\mathsf{T}}\dot{a}$ is a relative translation velocity between the two observers, while $q$ is the axial vector of the skew-symmetric second-rank tensor $Q^{\mathsf{T}}\dot{Q}$.
The vector $\nu:=\tilde{\nu}(x,t)\in\tilde{\mathbb{R}^{3}}$ collects degrees of freedom that are internal (thus relative) in time to the material element associated with $x\in\mathcal{B}$ in the reference configuration. This relative character renders $\nu$ a Galilean invariant vector, meaning it is insensitive to relative rigid translations between two generic observers. When $\mathcal{O}$ and $\mathcal{O}^{\prime}$ rotate relatively, the value $\nu$ recorded by a first observer changes into $\nu^{\prime}=Q(t)\nu$ and the time rate $\dot{\nu}$ becomes $\dot{\nu}^{\prime}=\dot{\nu}+\dot{Q}\nu$. By pulling back $\dot{\nu}^{\prime}$ into $\mathcal{O}$, we obtain a new vector $\dot{\nu}^{\diamond}:=Q^{\mathsf{T}}\dot{\nu}^{\prime}$ given by
$\dot{\nu}^{\diamond}=\dot{\nu}+q\times\nu\;,$
which we rewrite as
$\dot{\nu}^{\diamond}=\dot{\nu}+\mathcal{A}(\nu)q\;,$
where the linear operator $\mathcal{A}$ is given by $\mathcal{A}:=\mathcal{A}(\nu)=-\nu\times$.
3.2 External power and invariance requirement: consequences
A subset $\mathfrak{b}$ of $\mathcal{B}$ is said to be a part of $\mathcal{B}$ when it has non-vanishing volume and the same geometrical regularity required of $\mathcal{B}$.
External actions over $\mathfrak{b}$ are associated with $u$ and $\nu$; they are defined by the power that they perform in the time rates of the two fields considered. Such actions are subdivided, as usual, into bulk and contact families: those associated with $u$ are standard bulk forces and boundary traction, while those performing power in $\dot{\nu}$ are microstructural actions. The external power $\mathcal{P}^{ext}_{\mathfrak{b}}$ that they perform on any time rate of change in the body morphology is defined in Lagrangian representation to be
$\mathcal{P}_{\mathfrak{b}}^{ext}\left(\dot{y},\dot{\nu}\right):=\int_{\mathfrak{b}}\left(b^{\ddagger}\cdot\dot{y}+\beta^{\ddagger}\cdot\dot{\nu}\right)\mathrm{d}x+\int_{\partial\mathfrak{b}}\left(\mathfrak{t}_{\partial}\cdot\dot{y}+\tau_{\partial}\cdot\dot{\nu}\right)\mathrm{d}\mathcal{H}^{2}(x),$
where $\mathrm{d}\mathcal{H}^{2}(x)$ denotes the Hausdorff two-dimensional measure. The symbol $\partial$ as a subscript indicates that the contact actions defined over $\partial\mathfrak{b}$ depend on the boundary, in addition to $x$ and $t$. The bulk force $b^{\ddagger}$ is presumed to be the sum of inertial $b^{in}$ and non-inertial $b$ components, the former to be identified; $\beta^{\ddagger}$ also admits by assumption a similar decomposition: $\beta^{\ddagger}=\beta^{in}+\beta$.
We require that $\mathcal{P}_{\mathfrak{b}}^{ext}$ is invariant under isometry-based observer changes; so, we impose
$\mathcal{P}_{\mathfrak{b}}^{ext}\left(\dot{y},\dot{\nu}\right)=\mathcal{P}_{\mathfrak{b}}^{ext}\left(\dot{y}^{\diamond},\dot{\nu}^{\diamond}\right)$
for any choice of $\mathfrak{b}$, $\mathfrak{c}$, and $q$. It means that we consider $\mathcal{P}_{\mathfrak{b}}^{ext}$ to be objective. The arbitrariness of $\mathfrak{c}$ implies the standard integral balance of forces
$\int_{\mathfrak{b}}b^{{\ddagger}}\;\;\mathrm{d}x+\int_{\partial\mathfrak{b}}\mathfrak{t}_{\partial}\;\;\mathrm{d}\mathcal{H}^{2}(x)=0\;,$ (1)
while the one of $q$ a non-standard balance of couples
$\int_{\mathfrak{b}}\left(\left(y-y_{0}\right)\times b^{{\ddagger}}+\mathcal{A}^{\top}\beta^{{\ddagger}}\right)\mathrm{d}x+\int_{\partial\mathfrak{b}}\left(\left(y-y_{0}\right)\times\mathfrak{t}_{\partial}+\mathcal{A}^{\top}\tau_{\partial}\right)\mathrm{d}\mathcal{H}^{2}(x)=0\;.$ (2)
Equation (1) gives standard results. Precisely, if $\left|b^{{\ddagger}}\right|$ is bounded over $\mathcal{B}$ and $\mathfrak{t}_{\partial}$ depends continuously on $x$, Cauchy’s theorem on fluxes implies that $\mathfrak{t}_{\partial}$ depends on $\mathfrak{\partial\mathfrak{b}}$ only through the normal $\mathbf{n}$ to it in all points where $\mathbf{n}$ is well-defined; $\mathfrak{t}_{\partial}$ is also such that $\mathfrak{t}_{\partial}=\mathfrak{t}:=\tilde{\mathfrak{t}}\left(x,t,\mathbf{n}\right)=-\tilde{\mathfrak{t}}\left(x,t,-\mathbf{n}\right)$, and, as a function of $\mathbf{n}$, $\tilde{\mathfrak{t}}$ is homogeneous and additive, namely $\tilde{\mathfrak{t}}\left(x,t,\mathbf{n}\right)=P\left(x,t,\right)\mathbf{n}\left(x\right)$, with $P$ the first Piola-Kirchhoff stress.
Less standard are the consequences of the integral balance (2). Precisely, since $\mathcal{B}$ is bounded, we can choose the arbitrary point $y_{0}$ in a way such that the boundedness of $|b^{{\ddagger}}|$ implies the one of $|(y-y_{0})\times b^{{\ddagger}}|$. If $\left|\mathcal{A}^{\top}\beta^{{\ddagger}}\right|$ is also bounded and $\tau_{\partial}$ depends continuously on $x$, by exploiting once again Cauchy’s theorem, we realize that $\tau_{\partial}$ depends on $\mathfrak{\partial\mathfrak{b}}$ only through the normal $\mathbf{n}$; also, $\mathcal{A}^{\top}\left(\tilde{\tau}\left(x,t,\mathbf{n}\right)+\tilde{\tau}\left(x,t,-\mathbf{n}\right)\right)=0$ and, as a function of $\mathbf{n}$, $\tilde{\tau}$ is homogeneous and additive: $\tilde{\tau}\left(x,t,\mathbf{n}\right)=\mathcal{S}\left(x,t\right)\mathbf{n}\left(x\right)$. The second-rank tensor $\mathcal{S}$ is named a microstress.
Due to the arbitrariness of $\mathfrak{b}$, if both stress fields are in $C^{1}\left(\mathcal{B}\right)\cap C\left(\bar{\mathcal{B}}\right)$ and the bulk actions $x\longmapsto b$, $x\longmapsto\beta^{{\ddagger}}$ are continuous over $\mathcal{B}$, from (1) we get the standard point-wise balance of forces
$\mathrm{Div}P+b^{{\ddagger}}=0\;,$ (3)
while from (2) we get that there exists a field $(x,t)\longmapsto\mathbf{z}\left(x,t\right)\in\hat{\mathbb{R}}^{3}$ such that
$\mathrm{Div}\mathcal{S}+\beta^{{\ddagger}}-\mathbf{z}=0\;,\qquad\text{and}\qquad\mathrm{skw}(PF^{\top})=\frac{1}{2}\mathsf{e}\left(\mathcal{A}^{\top}\mathbf{z}+\left(\nabla\mathcal{A}^{\top}\right)\mathcal{S}\right)\;,$ (4)
where $\mathrm{skw}(\cdot)$ extract the skew-symmetric part of its argument; moreover,
$\mathcal{P}_{\mathfrak{b}}^{ext}\left(\dot{y},\dot{\nu}\right)=\int_{\mathfrak{b}}\left(P\cdot\dot{F}+\mathbf{z}\cdot\dot{\nu}+\mathcal{S}\cdot\dot{N}\right)\mathrm{d}x\;,$ (5)
where the right-hand side integral is called internal (or inner) power (see details in the case of general manifold-valued phase fields in [21] and [23]).
The inertial terms are identified by assuming that their power equals the negative of the kinetic energy time rate for any choice of the velocity fields. This approach is common. The kinetic energy of a generic complex body is presumed to be the sum of a standard quadratic expression $\frac{1}{2}\rho|\dot{y}|^{2}$, where $\rho$ is the mass density, plus a non-negative function $\kappa(\nu,\dot{\nu})$ such that $\kappa(\nu,0)=0$ and $\frac{\partial^{2}\kappa}{\partial\dot{\nu}\partial\dot{\nu}}\cdot\dot{\nu}\otimes\dot{\nu}\geq 0$ for generic phase fields [4] (for more general analyses on $\beta^{{\ddagger}}$ see [5], [21], [23]).
In the case of quasicrystals that we treat here with details, $\tilde{\nu}$ coincides with the so-called phason field; for such a material the possibility of peculiar inertia effects attributed to $\tilde{\nu}$ is questionable since only three sound-like branches are apparently recorded in experiments [33]. Consequently, even if only for later purposes, we leave apart the microstructural inertia and set $\kappa(\nu,\dot{\nu})=0$, so $\beta^{in}=0$ to within a powerless term that we also discard (for its presence in quasicrystals see [25] and [2]). The gross kinetic energy remains; the arbitrariness of $\dot{y}$ implies by continuity the identification $b^{in}=-\rho\ddot{y}=-\rho\ddot{u}$.
4 Constitutive structures
We write the Clausius-Duhem inequality in isothermal setting as
$\frac{\mathrm{d}}{\mathrm{d}t}\int_{\mathfrak{b}}\psi\;\;\mathrm{d}x-\int_{\mathfrak{b}}\left(P\cdot\dot{F}+\mathbf{z}\cdot\dot{\nu}+\mathcal{S}\cdot\dot{N}\right)\mathrm{d}x\leq 0\;,$
where $\psi$ is the free energy density. The inequality, a version of the second law of thermodynamics, is presumed to hold true for any choice of the time rates involved.
The arbitrariness of $\mathfrak{b}$ and the presumed continuity of the integrand imply a local dissipation inequality given by
$\dot{\psi}-P\cdot\dot{F}-\mathbf{z}\cdot\dot{\nu}-\mathcal{S}\cdot\dot{N}\leq 0\;.$
As constitutive functional dependence on state variables, we assume
$\psi=\tilde{\psi}(F,\nu,N)\;\qquad P=\tilde{P}(F,\nu,N)\;\qquad\mathcal{S}=\tilde{\mathcal{S}}(F,\nu,N)$
and
$\mathbf{z}=\tilde{\mathbf{z}}^{e}(F,\nu,N)+\tilde{\mathbf{z}}^{d}(F,\nu,N,\dot{\nu})\;,$
where, we recall, $\mathbf{z}^{d}=\tilde{\mathbf{z}}^{d}(F,\nu,N,\dot{\nu})$ is a dissipative component of the self-action $\mathbf{z}$, while $\mathbf{z}^{e}=\tilde{\mathbf{z}}^{e}(F,\nu,N)$ is a conservative component determined by the free energy $\psi$.
These choices imply that we consider dissipation to be restricted only at microstructural level. Thus, arbitrariness of the time rates involved in the inequality necessarily implies
$P=\frac{\partial\psi}{\partial F}\;,\quad\mathbf{z}^{e}=\frac{\partial\psi}{\partial\nu}\;,\quad\mathcal{S}=\frac{\partial\psi}{\partial N}\;,\quad\mathbf{z}^{d}\cdot\dot{\nu}\geq 0.$
The last inequality is compatible with
$\mathbf{z}=\frac{\partial\psi}{\partial\nu}+\tilde{\varsigma}(\dots)\dot{\nu}\;,$
where $\tilde{\varsigma}(\dots)$ is a positive-valued state function, here chosen to be a scalar $\varsigma>0$ only for the sake of simplicity. With $\nu$ referring to a length scale $\lambda$, both $\mathcal{S}$ and $\mathbf{z}$ are thus of order $\lambda$. Consequently, from equation (4), we obtain that $PF^{\top}$ is symmetric to within terms of the order $\lambda^{2}$.
$P$, $\mathcal{S}$, and $\mathbf{z}$ are values of fields defined over $\mathcal{B}$ in terms of space variables. They have counterparts defined over $\mathcal{B}_{c}:=\tilde{y}(\mathcal{B},t)$ and given by
$\sigma=\frac{1}{\mathrm{det}F}\frac{\partial\psi}{\partial F}F^{\top}\;,\quad\mathbf{z}_{c}=\frac{1}{\mathrm{det}F}\left(\frac{\partial\psi}{\partial\nu}+\tilde{\varsigma}(\dots)\dot{\nu}\right)\;,\quad\mathcal{S}_{c}=\frac{1}{\mathrm{det}F}\frac{\partial\psi}{\partial N}F^{\top},$
where $\sigma$ is the standard Cauchy stress, while $\mathbf{z}_{c}$ and $\mathcal{S}_{c}$ are respectively self-action and microstress in Eulerian representation.
In small strain regime, that is, we recall, when $|\nabla u|\ll 1$, we have the approximations
$\sigma\approx P\;,\quad\mathbf{z}_{c}\approx\mathbf{z}\;,\quad\mathcal{S}_{c}\approx\mathcal{S}\;.$
It is also possible to choose for the free energy density $\psi$ a quadratic form with respect to its entries, a choice otherwise forbidden at least with respect to $F$ because, as it is well-known, a quadratic dependence on $F$ would be incompatible with the natural request of objectivity for the energy density, that is a requirement of invariance under the action of $SO(3)$, hence with respect to relatively rotating observers.
By maintaining ourselves in the small strain setting, we thus choose a form of energy derived in [25] under symmetry requirements. It reads
$\psi=$ $\frac{1}{2}\lambda\left(\mathrm{sym}\nabla u\cdot I\right)^{2}+\mu\,\mathrm{sym}\nabla u\cdot\mathrm{sym}\nabla u$
$+\frac{1}{2}k_{1}\left(N\cdot I\right)^{2}+k_{2}\,\mathrm{sym}N\cdot\mathrm{sym}N+k_{2}^{\prime}\,\mathrm{skw}N\cdot\mathrm{skw}N$
$+k_{3}\left(\mathrm{sym}\nabla u\cdot I\right)\left(N\cdot I\right)+k_{3}^{\prime}\,\mathrm{sym}N\cdot\mathrm{sym}\nabla u+\frac{1}{2}k_{0}\left|\nu\right|^{2}$
where $I$ is once again the unit tensor and $\mathrm{sym}(\cdot)$ extracts the symmetric component of its argument; $\lambda$ and $\mu$ are the Lamé constants, the other coefficients indicated by $k$ with various decorations are related to $\nu$. They must be such that the energy is non-negative definite, thus $\mu>0$, $k_{2}>0$, $02k_{2}+3k_{1}>$ 0, $k_{2}^{\prime}>0$, $k_{3}^{\prime}<2\sqrt{\mu k_{2}}$, $3k_{3}+k_{3}^{\prime}<\sqrt{(2\mu+3\lambda)(2k_{2}+3k_{1})}$, $k_{0}\geq 0$.
We thus have
$P\approx\boldsymbol{\sigma}=\lambda\left(\mathrm{tr}\,(\mathrm{sym}\nabla u)\right)I+2\mu\mathrm{sym}\nabla u+k_{3}\left(\mathrm{tr}\nabla\nu\right)I+k_{3}^{\prime}\mathrm{sym}\nabla\nu+\epsilon\nabla\dot{u}\;,$
$\mathbf{z}\approx\mathbf{z}_{c}=k_{0}\nu+\varsigma\dot{\nu}\;,$
$\!\!\begin{array}[]{ll}\mathcal{S}\approx\mathcal{S}_{c}=&k_{1}\left(\mathrm{tr}\nabla\nu\right)I+2k_{2}\mathrm{sym}\nabla\nu+2k_{2}^{\prime}\mathrm{skw}\nabla\nu+k_{3}\left(\mathrm{tr}(\mathrm{sym}\nabla u)\right)I+k_{3}^{\prime}\mathrm{sym}\nabla u\end{array}$
By setting $\xi=\lambda+\mu$, $\alpha=k_{3}+\frac{1}{2}k^{\prime}_{3}$, $\zeta=k_{2}+k^{\prime}_{2}$, $\gamma=k_{1}+k_{2}-k^{\prime}_{2}$, $\chi=\frac{1}{2}k^{\prime}_{3}\;,$ and choosing $\beta=b=0$, the local balances of momentum and microstructural actions reduce to
$\rho\ddot{u}=\mu\Delta u+(\lambda+\mu)\nabla\mathrm{div}\,u+\chi\Delta\nu+\alpha\nabla\mathrm{div}\,\nu\;,$ (6)
and a balance of microstructural actions given by
$c\dot{\nu}+k_{0}\nu=\chi\Delta u+\alpha\nabla\mathrm{div}\,u+\zeta\Delta\nu+\gamma\nabla\mathrm{div}\,\nu\;.$ (7)
The analyses reported in what follows refer to them.
5 Waves with time-dependent amplitude
Assume
$u(x,t)=\mathbf{u}(t)e^{i\mathbf{k}\cdot\mathbf{x}}\quad\text{and}\quad\nu(x,t)=\mathbf{\nu}(t)e^{i\mathbf{k}\cdot\mathbf{x}}\;,$ (8)
where, we repeat, $\mathbf{k}=k\mathbf{n}$ is a constant vector, the wave amplitudes $\mathbf{u},\mathbf{\nu}\in\mathbb{R}^{3}$ depend on time, and $\mathbf{x}=x-x_{0}$, with $x_{0}$ an arbitrary point.
We consider various possible choices for the phason self-action in order to evaluate the effects of the action itself.
5.1 In the absence of a self-action
If $\mathbf{z}=0$, by inserting expressions (8) into the balance equations, we get
$\ddot{\mathbf{u}}=-k^{2}\bigg{(}\frac{\mu}{\rho}I+\frac{\lambda+\mu}{\rho}(\mathbf{n}\otimes\mathbf{n})\bigg{)}\mathbf{u}-\frac{k^{2}}{\rho}\bigg{(}\chi I+\alpha(\mathbf{n}\otimes\mathbf{n})\bigg{)}\mathbf{\nu}$ (9)
and
$\bigg{(}\zeta I+\gamma(\mathbf{n}\otimes\mathbf{n})\bigg{)}\mathbf{\nu}=-\bigg{(}\chi I+\alpha(\mathbf{n}\otimes\mathbf{n})\bigg{)}\mathbf{u}\;.$ (10)
By substituting (10) into (9) we get
$\ddot{\mathbf{u}}=\mathbb{K}\mathbf{u}\;,$ (11)
where
$\mathbb{K}:=-\frac{k^{2}}{\rho}\bigg{[}\left(\mu-\frac{\chi^{2}}{\zeta}\right)I+\left(\lambda+\mu-\frac{(\alpha+\chi)^{2}}{\zeta+\gamma}+\frac{\chi^{2}}{\zeta}\right)\mathbf{n}\otimes\mathbf{n}\bigg{]}\;.$ (12)
Thus, by defining
$\mathbf{q}:=\begin{bmatrix}\mathbf{u}\\ \mathbf{v}\end{bmatrix}\in\mathbb{R}^{6}\;,\qquad A:=\begin{bmatrix}\mathbf{0}&&&&I\\ \mathbb{K}&&&&\mathbf{0}\end{bmatrix}\in\mathbb{R}^{6\times 6}\;,$ (13)
where $\mathbf{v}:=\dot{\mathbf{u}}$, we rewrite (11) as
$\dot{\mathbf{q}}=A\mathbf{q}\;.$ (14)
We compute
$AA^{\top}=\begin{bmatrix}\mathbf{0}&&&&I\\ \mathbb{K}&&&&\mathbf{0}\end{bmatrix}\begin{bmatrix}\mathbf{0}&&&&\mathbb{K}^{\top}\\ I&&&&\mathbf{0}\end{bmatrix}=\begin{bmatrix}I&&&&\mathbf{0}\\ \mathbf{0}&&&&\mathbb{K}\mathbb{K}^{\top}\end{bmatrix}\;,$
while
$A^{\top}A=\begin{bmatrix}\mathbf{0}&&&&\mathbb{K}^{\top}\\ I&&&&\mathbf{0}\end{bmatrix}\begin{bmatrix}\mathbf{0}&&&&I\\ \mathbb{K}&&&&\mathbf{0}\end{bmatrix}=\begin{bmatrix}\mathbb{K}^{\top}\mathbb{K}&&&&\mathbf{0}\\ \mathbf{0}&&&&I\end{bmatrix}\;;$
therefore, $A$ is non-normal (normality being defined by the equality $AA^{\top}=A^{\top}A$ when $A\in\mathbb{R}^{n\times n}$ or by $AA^{\ast}=A^{\ast}A$, when $A\in\mathbb{C}^{n\times n}$).
Nonetheless, even if $A$ is non-normal, the stability of the system is completely described by its eigenvalues since
$A^{2}=\begin{bmatrix}\mathbf{0}&&&&I\\ \mathbb{K}&&&&\mathbf{0}\end{bmatrix}\begin{bmatrix}\mathbf{0}&&&&I\\ \mathbb{K}&&&&\mathbf{0}\end{bmatrix}=\begin{bmatrix}\mathbb{K}&&&&\mathbf{0}\\ \mathbf{0}&&&&\mathbb{K}\end{bmatrix}\;.$ (15)
Therefore, the spectrum of $A$ is given by the square roots of the eigenvalues of the symmetric matrix $\mathbb{K}$, namely
$\sigma(A)=\bigcup_{\lambda\in\sigma(K)}\pm\sqrt{\lambda}\;.$
Hence, the eigenvalues of $A$ are either real or purely imaginary. In particular the eigenvalues of $A$ are real, both positive and negative, when $\mathbb{K}$ is positive definite; imaginary when $\mathbb{K}$ is negative definite and real and imaginary when $\mathbb{K}$ is indefinite. Moreover, the spectrum of $A$ is stable under (symmetric) perturbations in $\mathbb{K}$ thanks to Weyl’s Theorem [16], and thus the eigenvalues provide reliable information on the stability of the system.
5.2 In the presence of a conservative self-action
Le us consider $\mathbf{z}=k_{0}\nu$. The balance equations still reduce to a system of the form
$\dot{\mathbf{q}}=A\mathbf{q}\;,$ (16)
with
$\mathbf{q}:=\begin{bmatrix}\mathbf{u}\\ \mathbf{v}\end{bmatrix}\quad\text{and}\quad A:=\begin{bmatrix}\mathbf{0}&&&&I\\ \mathbb{K}&&&&\mathbf{0}\end{bmatrix}\in\mathbb{R}^{6\times 6}\;,$ (17)
but now
$\mathbb{K}:=-\frac{k^{2}}{\rho}\bigg{[}\left(\mu-\frac{\chi^{2}k^{2}}{k_{0}+\zeta k^{2}}\right)I+\bigg{(}\lambda+\mu-k^{2}\left(\frac{(\alpha+\chi)^{2}}{k_{0}+k^{2}(\zeta+\gamma)}-\frac{\chi^{2}}{k_{0}+\zeta k^{2}}\right)\bigg{)}\mathbf{n}\otimes\mathbf{n}\bigg{]}\;.$ (18)
As in the previous case, all the eigenvalues of $A$ are either real or purely imaginary. In order to analyze the system stability it suffices to compute the spectrum of the symmetric matrix $\mathbb{K}$: if $\mathbb{K}$ is either indefinite or negative definite, the system is unstable. Once again, the eigenvalues of $A$ are stable under small symmetric perturbations on $\mathbb{K}$.
5.3 A purely dissipative self-action
Assume that $\mathbf{z}$ is endowed only with a dissipative component driving diffusion, namely it is of the form $\mathbf{z}=\varsigma\dot{\nu}$, with $\varsigma$ a positive constant. From the balance equations we obtain
$\ddot{\mathbf{u}}=-k^{2}\bigg{(}\frac{\mu}{\rho}I+\frac{\lambda+\mu}{\rho}(\mathbf{n}\otimes\mathbf{n}\mathbf{)}\bigg{)}\mathbf{u}-\frac{k^{2}}{\rho}\bigg{(}\chi I+\alpha(\mathbf{n}\otimes\mathbf{n}\mathbf{)}\bigg{)}\mathbf{\nu}$ (19)
and
$\dot{\mathbf{\nu}}=-\frac{k^{2}}{\varsigma}\bigg{(}\chi I+\alpha(\mathbf{n}\otimes\mathbf{n}\mathbf{)}\bigg{)}\mathbf{u}-k^{2}\bigg{(}\frac{\zeta}{\varsigma}I+\frac{\gamma}{\varsigma}(\mathbf{n}\otimes\mathbf{n}\mathbf{)}\bigg{)}\mathbf{\nu}\;,$ (20)
which can be rewritten as
$\rho\ddot{\mathbf{u}}=\mathbb{K}_{1}\mathbf{u}+\frac{1}{\rho}\mathbb{K}_{2}\mathbf{\nu}\;,\qquad\dot{\mathbf{\nu}}=\frac{1}{\varsigma}\mathbb{K}_{2}\mathbf{u}+\mathbb{K}_{3}\mathbf{\nu}\;,$ (21)
where
$\mathbb{K}_{1}:=-k^{2}\bigg{(}\frac{\mu}{\rho}I+\frac{\lambda+\mu}{\rho}(\mathbf{n}\otimes\mathbf{n}\mathbf{)}\bigg{)}\;,$ (22)
$\mathbb{K}_{2}:=-k^{2}\bigg{(}\chi I+\alpha(\mathbf{n}\otimes\mathbf{n}\mathbf{)}\bigg{)}\;,$ (23)
$\mathbb{K}_{3}:=-k^{2}\bigg{(}\frac{\zeta}{\varsigma}I+\frac{\gamma}{\varsigma}(\mathbf{n}\otimes\mathbf{n}\mathbf{)}\bigg{)}\;.$ (24)
By setting
$\mathbf{q}:=\begin{bmatrix}\mathbf{u}\\ \mathbf{v}\\ \mathbf{\nu}\end{bmatrix}\in\mathbb{R}^{9}\;,\quad A:=\begin{bmatrix}\mathbf{0}&I&\mathbf{0}\\ \mathbb{K}_{1}&\mathbf{0}&\frac{1}{\rho}\mathbb{K}_{2}\\ \frac{1}{\varsigma}\mathbb{K}_{2}&\mathbf{0}&\mathbb{K}_{3}\end{bmatrix}\in\mathbb{R}^{9\times 9}\;,$ (25)
where, as above, $\mathbf{v}:=\dot{\mathbf{u}}$, we rewrite system (21) in the standard form
$\dot{\mathbf{q}}=A\mathbf{q}\;.$ (26)
The matrix $A$ is non-normal because
$AA^{\top}=\begin{bmatrix}\mathbf{0}&I&\mathbf{0}\\ \mathbb{K}_{1}&\mathbf{0}&\frac{1}{\rho}\mathbb{K}_{2}\\ \frac{1}{\varsigma}\mathbb{K}_{2}&\mathbf{0}&\mathbb{K}_{3}\end{bmatrix}\begin{bmatrix}\mathbf{0}&\mathbb{K}_{1}^{\top}&\frac{1}{\varsigma}\mathbb{K}_{2}^{\top}\\ I&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\frac{1}{\rho}\mathbb{K}_{2}^{\top}&\mathbb{K}_{3}^{\top}\end{bmatrix}=\begin{bmatrix}I&0&0\\ 0&\mathbb{K}_{1}\mathbb{K}_{1}^{\top}+\frac{1}{\rho^{2}}\mathbb{K}_{2}\mathbb{K}_{2}^{\top}&\frac{1}{\varsigma}\mathbb{K}_{1}\mathbb{K}_{2}^{\top}+\frac{1}{\rho}\mathbb{K}_{2}\mathbb{K}_{3}^{\top}\\ 0&\frac{1}{\varsigma}\mathbb{K}_{2}\mathbb{K}_{1}^{\top}+\frac{1}{\rho}\mathbb{K}_{3}\mathbb{K}_{2}^{\top}&\frac{1}{\varsigma^{2}}\mathbb{K}_{2}\mathbb{K}_{2}^{\top}+\mathbb{K}_{3}\mathbb{K}_{3}^{\top}\end{bmatrix}$
and
$A^{\top}A=\begin{bmatrix}\mathbf{0}&\mathbb{K}_{1}^{\top}&\frac{1}{\varsigma}\mathbb{K}_{2}^{\top}\\ I&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\frac{1}{\rho}\mathbb{K}_{2}^{\top}&\mathbb{K}_{3}^{\top}\end{bmatrix}\begin{bmatrix}\mathbf{0}&I&\mathbf{0}\\ \mathbb{K}_{1}&\mathbf{0}&\frac{1}{\rho}\mathbb{K}_{2}\\ \frac{1}{\varsigma}\mathbb{K}_{2}&\mathbf{0}&\mathbb{K}_{3}\end{bmatrix}=\begin{bmatrix}\mathbb{K}_{1}^{\top}\mathbb{K}_{1}+\frac{1}{\varsigma^{2}}\mathbb{K}_{2}^{\top}\mathbb{K}_{2}&0&\frac{1}{\rho}\mathbb{K}_{1}^{\top}\mathbb{K}_{2}+\frac{1}{\varsigma}\mathbb{K}_{2}^{\top}\mathbb{K}_{3}\\ 0&I&0\\ \frac{1}{\rho}\mathbb{K}_{2}^{\top}\mathbb{K}_{1}+\frac{1}{\varsigma}\mathbb{K}_{3}^{\top}\mathbb{K}_{2}&0&\frac{1}{\rho^{2}}\mathbb{K}_{2}^{\top}\mathbb{K}_{2}+\mathbb{K}_{3}^{\top}\mathbb{K}_{3}\end{bmatrix}\;.$
At variance with the previous two cases, the non-normality in $A$ is now genuine, and we can no longer be sure that small perturbations on $\mathbb{K}_{i}$ will result in small changes in the spectrum of $A$. In other words, the spectrum of $A$ could be highly sensitive to small changes in the parameters, making stability predictions based on the eigenvalues alone unreliable, especially in the presence of uncertainties or rounding errors in computed quantities. To assess stability we need to have a look at the pseudospectrum of $A$.
5.4 Simultaneous presence of conservative and dissipative components in the self-action
Finally, we consider $\mathbf{z}=k_{0}\nu+\varsigma\dot{\nu}$, so that, with previous notations, the matrix $A$ in the system
$\dot{\mathbf{q}}=A\mathbf{q}\;,$ (27)
becomes
$A:=\begin{bmatrix}\mathbf{0}&I&\mathbf{0}\\ \mathbb{K}_{1}&\mathbf{0}&\frac{1}{\rho}\mathbb{K}_{2}\\ \frac{1}{\varsigma}\mathbb{K}_{2}&\mathbf{0}&\mathbb{K}_{3}\end{bmatrix}\;,$ (28)
with
$\mathbb{K}_{1}:=-k^{2}\bigg{(}\frac{\mu}{\rho}I+\frac{\lambda+\mu}{\rho}(\mathbf{n}\otimes\mathbf{n}\mathbf{)}\bigg{)}\;,$ (29)
$\mathbb{K}_{2}:=-k^{2}\bigg{(}\chi I+\alpha(\mathbf{n}\otimes\mathbf{n}\mathbf{)}\bigg{)}\;,$ (30)
$\mathbb{K}_{3}:=-k^{2}\bigg{(}\frac{\zeta+\frac{k_{0}}{k^{2}}}{c}I+\frac{\gamma}{c}(\mathbf{n}\otimes\mathbf{n}\mathbf{)}\bigg{)}\;.$ (31)
Similar considerations to those made in the previous case apply here.
6 Pseudospectrum: general properties
6.1 Reasons for calling upon the pseudospectrum
For a system of the form (27), information on the asymptotic behavior of solutions are given by the spectrum of $A$. Specifically, as it is well-known, when all the eigenvalues of $A$ have a negative or null real part the system is said to be stable. In contrast, an amplified wave to infinity is associated with an eigenvalue of positive real part; in this case the system is said to be unstable. A parametric analysis may reveal possible transitions from stability to instability and vice versa.
The analysis along these lines is straightforward when $A$ is a normal matrix. The eigenvalues indicate characteristic times of exponentially relaxing modes whose linear combination is a solution to the considered system. When small perturbations are accounted for, small is the variation of the $A$ spectrum. At variance, when $A$ is non-normal, as in cases we tackle here, small perturbations can have significant effects that may induce instability. More specifically, for matrices far from being normal the standard spectral analysis shows possible failures in the evaluation of resonance, asymptotic behavior, diagonalization, geometric characterization [36]. Moreover, little or no information on the transient behavior and rate of convergence to steady-state can be gleaned in the highly non-normal case from the eigenvalues alone [1].
A way to manage such aspects is to consider an extended notion of spectrum, what is called a $\varepsilon$-pseudospectrum [36], [13], [26], [34].
If there exists a small $\varepsilon$ such that the $\varepsilon$-pseudospectrum trespasses in the half plane with positive real part, the system has to be treated as unstable.
6.2 Definitions and a few properties
Let $A$ be a $n\times n$ matrix with complex entries; in short, $A\in\mathbb{C}^{n\times n}$. The set $\sigma(A)$ of $A$ eigenvalues contains all $\lambda\in\mathbb{C}$ satisfying $\det(A-\lambda I)=0$, with $I$ the $n\times n$ unit matrix. The singular character of a matrix is however perturbation sensitive: if $A\in\mathbb{C}^{n\times n}$ is singular, there exists always $E\in\mathbb{C}^{n\times n}$, with $\|E\|_{2}\ll 1$ such that the matrix $A+E$ is non-singular. So, instead of asking whether $\lambda I-A$ is singular, we may look at the norm of its inverse, checking whether the value is sufficiently large. Of course, choosing the norm of its inverse matters; here we refer to the matrix norm induced by the Euclidean one.
We have three equivalent definitions of $\varepsilon$-pseudospectrum [36].
Definition 1 (First version)
Given $A\in\mathbb{C}^{n\times n}$, take an arbitrary $\varepsilon>0$. The $\varepsilon$-pseudospectrum of $A$, indicated by $\sigma_{\varepsilon}(A)$, is the set of $\zeta\in\mathbb{C}$ such that
$\|(\zeta I-A)^{-1}\|_{2}>\frac{1}{\varepsilon}\;.$ (32)
Here we use the convention $\|(\zeta I-A)^{-1}\|_{2}=\infty$ if $\zeta I-A$ is singular.
Definition 2 (Second version)
The $\varepsilon$-pseudospectrum $\sigma_{\varepsilon}(A)$ is the set of complex numbers $\zeta\in\mathbb{C}$ such that
$\zeta\in\sigma(A+E),$ (33)
for some $E\in\mathbb{C}^{n\times n}$ with $\|E\|_{2}<\varepsilon$.
Definition 3 (Third version)
The $\varepsilon$-pseudospectrum $\sigma_{\varepsilon}(A)$ is the set of complex number $\zeta\in\mathbb{C}$ such that
$\|(\zeta I-A)\mathbf{v}\|_{2}<\epsilon$ (34)
for some $\mathbf{v}\in\mathbb{C}^{n\times n}$, with $\|\mathbf{v}\|_{2}=1$.
The idea – we repeat – is thus to compute the spectra of matrices differing from $A$ by a perturbation given by matrices in a ball of radius $\varepsilon$, evaluating in this way how far the spectrum of $A$ can go under perturbations.
Assume that $A$ is diagonalizable and is not necessarily normal. Indicate by $V$ the matrix of eigenvectors $\mathbf{v}$ of $A$, namely
$V=\begin{bmatrix}\mathbf{v}_{1}&\dots&\mathbf{v}_{n}\end{bmatrix}\in\mathbb{C}^{n\times n}\;,$
where each column $\mathbf{v}_{j}$ is normalized by $\|\mathbf{v}_{j}\|_{2}=1$. The number
$\kappa_{2}(V):=\|V\|_{2}\|V^{-1}\|_{2}=\frac{s_{\max}(V)}{s_{\min}(V)}\;,$ (35)
where $s_{\max}(V)$ and $s_{\min}(V)$ are the maximum and minimim singular values of $V$, is called the spectral condition number of $V$. In general, $1\leq\kappa_{2}(V)<\infty$; specifically, if $A$ is a normal matrix, we have $\kappa_{2}(V)=1$. It is known that if $A$ is far from normal, then $\kappa_{2}(V)$ will be large. While the converse is not true in general, an ill-conditioned $V$ often signals that at least one of the eigenvalues of $A$ will be sensitive to small perturbations. Hence, one can use the condition number of $V$, $\kappa_{2}(V)\gg 1$, as a measure of non-normality, albeit with caution.
Another tool for detecting possible eigenvalue instabilities is the computation of the $\varepsilon$-pseudospectrum for suitable values of $\varepsilon$. Given $\varepsilon>0$, the pseudospectrum of $A\in\mathbb{C}^{n\times n}$, namely $\sigma_{\varepsilon}(A)$, referred to the Euclidan norm, satisfies what follows (see [36] for the pertinent proofs):
  1. 1.
    $\sigma_{\varepsilon}(A)$ is an open bounded non-empty set composed of at most $n$ connected subsets, each containing one or more eigenvalues of $A$.
  2. 2.
    $\sigma_{\varepsilon}(A^{*})=\overline{\sigma_{\varepsilon}(A)}$, the conjugate of the pseudospectrum of $A$.
  3. 3.
    $\sigma_{\varepsilon}(A_{1}\oplus A_{2})=\sigma_{\varepsilon}(A_{1})\cup\sigma_{\varepsilon}(A_{2})$, where
    $A_{1}\oplus A_{2}=\begin{bmatrix}A_{1}&0\\ 0&A_{2}\end{bmatrix}\;.$
  4. 4.
    $\forall c\in\mathbb{C}$, $\sigma_{\varepsilon}(A+cI)=c+\sigma_{\varepsilon}(A)$.
  5. 5.
    $\forall c\in\mathbb{C}$, $c\not=0$, we get
    $\sigma_{|c|\varepsilon}(cA)=c\sigma_{\varepsilon}(A)\;.$
    In particular, for $c=2$ we have
    $\sigma_{\varepsilon}(2A)=2\sigma_{\frac{\varepsilon}{2}}(A)\;.$
Since the matrices involved in our analysis are real, it seems more appropriate to perturb the system matrix $A\in\mathbb{R}^{N\times N}$ with real perturbation matrices $E\in\mathbb{R}^{N\times N}$.
Definition 4
The real structured $\varepsilon$-pseudospectrum of a matrix $A\in\mathbb{R}^{N\times N}$ is defined as
$\sigma_{\varepsilon}^{\mathbb{R}}(A)=\bigcup_{\begin{subarray}{c}E\in\mathbb{R}^{N\times N}\\ \|E\|<\varepsilon\end{subarray}}\sigma(A+E)\;.$ (36)
The real $\varepsilon$-pseudospectrum is of course a subset of the complex $\varepsilon$-pseudospectrum, $\sigma_{\varepsilon}^{\mathbb{R}}(A)\subset\sigma_{\varepsilon}(A)$.
6.3 Structured pseudospectra
The matrices (25) and (28) have a block-type structure, namely
$A=\begin{bmatrix}\mathbf{0}&&&I&&&\mathbf{0}\\ A_{1}&&&\mathbf{0}&&&A_{2}\\ A_{3}&&&\mathbf{0}&&&A_{4}\end{bmatrix}\in\mathbb{R}^{9\times 9}\;.$ (37)
In view of this structure, it may be appropriate to perturb only the blocks that are different from the identity and the null matrix. The notion of structured pseudospectrum enters thus into play [29], [35], [36].
Definition 5
The structured $\varepsilon$-pseudospectrum of a matrix $A\in\mathbb{R}^{N\times N}$ of the form (37) is defined as
$\sigma_{\varepsilon}^{Str}(A)=\bigcup_{\|E\|<\varepsilon}\sigma(A+E)\;,$ (38)
with
$E=\begin{bmatrix}\mathbf{0}&&&\mathbf{0}&&&\mathbf{0}\\ E_{1}&&&\mathbf{0}&&&E_{2}\\ E_{3}&&&\mathbf{0}&&&E_{4}\end{bmatrix}\in\mathbb{R}^{9\times 9}\;.$ (39)
The complex pseudospectrum is computed adopting the close relation between the eigenvalues of the perturbed matrices and the resolvent norm. In the case of real perturbations, such a relation does not exist. Thus we proceed as follows: For any $\zeta\not\in\sigma(A)$, denote the resolvent by $\mathbf{R}(\zeta)=(\zeta I-A)^{-1}$. The 2-norm of the smallest real matrix $E$ that makes $\zeta$ an eigenvalue of $A+E$ is given by
$p(A,\zeta)\equiv\left(\inf_{\gamma\in(0,1]}s_{2}\left(\begin{bmatrix}\mathrm{Re}\,\mathbf{R}(\zeta)&&&-\gamma\mathrm{Im}\,\mathbf{R}(\zeta)\\ \gamma^{-1}\mathrm{Im}\,\mathbf{R}(\zeta)&&&\mathrm{Re}\,\mathbf{R}(\zeta)\end{bmatrix}\right)\right)^{-1}\;,$ (40)
where $s_{2}$ is the second-largest singular value. Thus, the 2-norm real structured $\varepsilon$-pseudospectrum of $A\in\mathbb{R}^{N\times N}$ is given by [36, Theorem 50.1]
$\sigma_{\varepsilon}^{\mathbb{R}}(A)=\{\zeta\in\mathbb{C}\;|\;p(A,\zeta)<\varepsilon\}\;.$ (41)
In the special case of quasicrystals treated here, however, the pseudospectrum analyses do not differ in essence from their structured counterparts. Figure 1 exemplifies such an aspect. Of course, it will not be always necessarily so.
7 Pseudospectrum in the quasicrystal dynamics: implications
We refer explicit computations to quasicrystals, as already anticipated. The software used is the one developed by Wright [40]. Data considered are as follows:
$\lambda=85\,\mathrm{GPa}\;,\quad\mu=65\,\mathrm{GPa}\;,\quad\zeta=0.044\,\mathrm{GPa}\;,\quad\gamma=0.0198\,\mathrm{GPa}\;.$
Mass density is $\rho=5100\,\mathrm{kg\,m^{-3}}$ [39], [19]. Values of the coupling parameters $\chi$ and $\alpha$ as well as those of $k_{0}$ and $\varsigma$, which enter the self-action, are unknown. However, we know a theoretical range for the coupling parameter $\chi$, which is $[0,1.5]\,\mathrm{GPa}$, as established in [28]. We also take the friction coefficient $\varsigma$ of the form $\varsigma=e^{\mathfrak{N}}\,\mathrm{Pa\cdot s/m^{2}}$, with $\mathfrak{N}\in[15,19]$.
The focus is on solutions of the form
$u=\mathrm{Re}\{\mathbf{u}(t)e^{i(\mathbf{k}\cdot\mathbf{x})}\}\;,\quad\nu=\mathrm{Re}\{\mathbf{\nu}(t)e^{i(\mathbf{k}\cdot\mathbf{x})}\}\;,$
where the time-dependent amplitudes $\mathbf{u}$ and $\mathbf{\nu}$ are complex.
In the following analysis of the pseudospectrum, perturbations $\varepsilon$ are of the order of Pa ($\varepsilon\approx 10^{-9}\chi$).
7.1 No phason self-action
Here we consider the matrix $A$ in (13) with $\mathbb{K}$ given by (12). Table 1 lists the eigenvalues of $A$ as $\chi$ varies and $\alpha=0$ GPa.
For $\chi<1.6912$ all the eigenvalues of $A$ are purely imaginary: the system undergoes infinitely many oscillations without damping.
When $\chi$ grows, the eigenvalues of $A$ that correspond to transversal waves migrate towards the real axis. Also, for $\chi\geq 1.6912$ GPa two coincident eigenvalues become real and positive, meaning that for such values of the parameters $\mathbb{K}$ is indefinite. Indeed, for $\chi=1.6912$ GPa and $\alpha=0$ GPa we have
[…]
Thus, for $\chi\geq 1.6912$ GPa instability occurs. The limit is in agreement with what has been theoretically foreseen in [28].
Table 1: No phason self-action: spectrum of $A$ for some values of $\chi$.
$\alpha=0$ GPa
$\chi$ $[\mathrm{GPa}]$ $0.05$ $0.5$ $1.6$ $1.6912$ $1.7$
$\sigma(A)$ $[\mathrm{rad/s}]$ $6492.24\,i$ $6433.39\,i$ $5855.69\,i$ $5776.39\,i$ $5768.44\,i$
$6492.24\,i$ $-6433.39\,i$ $-5855.69\,i$ $-5776.39\,i$ $-5768.44\,i$
$3568.47\,i$ $3410.43\,i$ $1156.24\,i$ ${\color[rgb]{1,0,0}+}26.49$ ${\color[rgb]{1,0,0}+}365.64$
$-3568.47\,i$ $-3410.43\,i$ $-1156.24\,i$ $-26.49$ $-365.64$
$3568.47\,i$ $3410.43\,i$ $1156.24\,i$ ${\color[rgb]{1,0,0}+}26.49$ ${\color[rgb]{1,0,0}+}365.64$
$-3568.47\,i$ $-3410.43\,i$ $-1156.24\,i$ $-26.49$ $-365.64$
7.2 Phason self-action with purely conservative character
When $\mathbf{z}=k_{0}\nu$ we refer to the matrix $A$ given by (17) for which $\mathbb{K}$ is given by formula (18).
[…]
Table 2: Conservative self-action: spectrum of $A$ for some reasonable values of $k_{0}$.
$\chi=0.1$ GPa, $\alpha=0$ GPa
$k_{0}$ $[\mathrm{GPa/m^{2}}]$ $0.001$ $0.01$ $0.1$ $1$
$\sigma(A)$ $[\mathrm{rad/s}]$ $6492.81\,i$ $6492.81\,i$ $6492.82\,i$ $6492.83\,i$
$-6492.81\,i$ $-6492.81\,i$ $-6492.82\,i$ $-6492.83\,i$
$3569.97\,i$ $3569.98\,i$ $3570.01\,i$ $3570.03\,i$
$-3569.97\,i$ $-3569.98\,i$ $-3570.01\,i$ $-3570.03\,i$
$3569.97\,i$ $3569.98\,i$ $3570.01\,i$ $3570.03\,i$
$-3569.97\,i$ $-3569.98\,i$ $-3570.01\,i$ $-3570.03\,i$
However, as $k_{0}$ increases, the stability range for the coupling parameter $\chi$ increases, meaning that the system becomes unstable for higher values of $\chi$. Specifically, for $k_{0}=0.01$ $\mathrm{GPa/m^{2}}$, instability occurs when $\chi\geq 1.88$ GPa; $k_{0}=0.1$ $\mathrm{GPa/m^{2}}$ corresponds to an unstable state for $\chi\geq 3.06$ GPa; the value $k_{0}=1$ $\mathrm{GPa/m^{2}}$ is a threshold towards instability for $\chi\geq 8.24$ GPa. In other words, the microstructural (phason in this case) self-action stabilizes the material.
7.3 Microstructural self-action with purely dissipative character
When $\mathbf{z}=\varsigma\dot{\nu}$, we consider the matrix in formula (25) with $\mathbb{K}_{i}$, $i=1,2,3$ given by formulas (22), (23), and (24).
In Figure 1 we compare complex and real $\varepsilon$-pseudospectra of $A$, the latter being a subset of the former, for different values of $\varepsilon$ in Pa, whose 10-based logarithm is displayed in the colorbar on the right of every snapshot. The material parameters are fixed to $\chi=0.5$ GPa and $\varsigma=0.178\,\mathrm{GPa\cdot s/m^{2}}$ ($\mathfrak{N}=19$). With these values, the departure from normality of $A$ given by the condition number $\kappa_{2}(V)$ is $\kappa(V)=6.49\cdot 10^{3}$.
(a) $\sigma_{\varepsilon}(A)$
(b) $\sigma_{\varepsilon}^{\mathbb{R}}(A)$
(c) $\sigma_{\varepsilon}^{Str}(A)$
Figure 1: Comparison between complex, real and structured $\varepsilon$-pseudospectrum of the matrix $A$ in (25) for $\chi=0.5$ GPa and $\varsigma=0.178\,\mathrm{GPa\cdot s/m^{2}}$. The figure on the right is made by the eigenvalues of 200 perturbation matrices of the form (39). In the figure (c) there is a zoom of the eigenvalue clouds: their values are, indeed, very close with each other.
Table 3 lists the eigenvalues of $A$ for $\alpha=0$ GPa, $\varsigma=0.178\,\mathrm{GPa\cdot s/m^{2}}$ as $\chi$ varies. When $\chi=1.7$ GPa the last two eigenvalues have positive real part: instability occurs. We find the same result by looking at the pseudospectrum in Figure 2 as $\chi$ varies, for $\varsigma=0.178\,\mathrm{GPa\cdot s/m^{2}}$, that is $\mathfrak{N}=19$, and $\alpha=0$ GPa. Once again instability is reached for $\chi>\sim 1.6$, so we recover numerically and from another path the theoretical result in [28] and we conclude that in this case the eigenvalues provide reliable information on the transition to instability.
Table 3: Dissipative self-action: spectrum of $A$ for some values of $\chi$.
$\alpha=0$ GPa, $\mathfrak{N}=19$
$\chi$ $[\mathrm{GPa}]$ $0.05$ $0.5$ $1.6$ $1.7$
$\sigma(A)$ $[\mathrm{rad/s}]$ $-3\cdot 10^{-5}+6493\,i$ $-3\cdot 10^{-3}+6493\,i$ $-0.03+6493\,i$ $-0.04+6493\,i$
$-3\cdot 10^{-5}-6493\,i$ $-3\cdot 10^{-3}-6493\,i$ $-0.03-6493\,i$ $-0.04-6493\,i$
$-0.36$ $-0.35$ $-0.29$ $-0.28$
$-1\cdot 10^{-4}+3570\,i$ $-0.01+3570\,i$ $-0.11+3570\,i$ $-0.12+3570\,i$
$-1\cdot 10^{-4}-3570\,i$ $-0.01-3570\,i$ $-0.11-3570\,i$ $-0.12-3570\,i$
$-1\cdot 10^{-4}+3570\,i$ $-0.01-3570\,i$ $-0.11+3570\,i$ $-0.12+3570\,i$
$-1\cdot 10^{-4}-3570\,i$ $-0.01+3570\,i$ $-0.11-3570\,i$ $-0.12-3570\,i$
$-0.25$ $-0.22$ $-0.03$ +$0.003$
$-0.25$ $-0.22$ $-0.03$ +$0.003$
Figure 2: Dissipative self-action: real pseudospectrum in a neighborhood of the real eigenvalues of $A$ as $\chi$ varies.
Fixing $\chi=0.1$ GPa and $\alpha=0$ GPa, as $\mathfrak{N}$ varies we obtain results collected in Table 4 and Figure 3. Once again we show the pseudospectra next to the real eigenvalues of $A$, which are the ones that become unstable when $\chi>1.6$ GPa. As $\mathfrak{N}$ grows, the imaginary part of the complex eigenvalues of $A$ diminishes. These eigenvalues correspond to damped oscillations and describe asymptotically stable states: the origin is an attracting point in the phase space. As $\varsigma$ grows, the damping decreases; also the real part of the eigenvalues grows as $\mathfrak{N}$ increases. Correspondingly, we have non-oscillating solutions: the displacement tends exponentially to zero from the initial value, without oscillations. Finally, the presence of a dissipative phason self-action does not stabilize the matter in the (numerical) admissible range for the coupling parameter: for $\mathfrak{N}$ varying from $15$ to $20$, instability always occurs for $\chi\geq 1.69$ GPa. In particular, when $\chi<1.69$ GPa, the spectral radius of $A$ goes to zero as $\mathfrak{N}$ increases.
Table 4: Dissipative self-action: spectrum of $A$ for some values of $\mathfrak{N}$.
$\alpha=0$ GPa, $\chi=0.1$ GPa
$\mathfrak{N}$ $18$ $19$ $20$
$\sigma(A)$ $[\mathrm{rad/s}]$ $-0.0004+6493\,i$ $-0.0001+6493\,i$ $-5\cdot 10^{-5}+6493\,i$
$-0.0004-6493\,i$ $-0.0001-6493\,i$ $-5\cdot 10^{-5}-6493\,i$
$-0.971$ $-0.357$ $-0.131$
$-0.001+3570\,i$ $-0.0004+3570\,i$ $-0.0002+3570\,i$
$-0.001-3570\,i$ $-0.0004-3570\,i$ $-0.0002-3570\,i$
$-0.001+3570\,i$ $-0.0004+3570\,i$ $-0.0002+3570\,i$
$-0.001-3570\,i$ $-0.0004-3570\,i$ $-0.0002-3570\,i$
$-0.668$ $-0.246$ $-0.09$
$-0.668$ $-0.246$ $-0.09$
Figure 3: Dissipative self-action: real pseudospectrum of $A$ as $\mathfrak{N}$ varies.
7.4 Simultaneous presence of conservative and dissipative components of z
When $\mathbf{z}=k_{0}\nu+\varsigma\dot{\nu}$, we consider the matrix in formula (28) with $\mathbb{K}_{i}$, $i=1,2,3$ given by formulas (29), (30), and (31).
Figure 4 and 5 show the complex $\varepsilon$-pseudospectrum of $A$ for $\chi=0.5$ GPa and $\chi=1.5$ GPa, respectively and for the same values of $\varepsilon$. Even if the matrix $A$ is genuinely non-normal, with $\kappa_{2}(V)=6.50\cdot 10^{3}$, its spectrum is stable under (generic) complex perturbations.
Figure 4: Complex $\varepsilon$-pseudospectrum of $A$ for $\chi=0.5$ GPa
Figure 5: Complex $\varepsilon$-pseudospectrum of $A$ for $\chi=0.5$ GPa
Table 5 collects eigenvalues of $A$ as $\chi$ varies, by fixing $\alpha=0$ GPa, $k_{0}=0.01$ $\mathrm{GPa/m^{2}}$, and $\varsigma=0.178\,\mathrm{GPa\cdot s/m^{2}}$, while Table 6 shows values of $\sigma(A)$ for $k_{0}=0.1$ $\mathrm{GPa/m^{2}}$, $\varsigma=0.178\,\mathrm{GPa\cdot s/m^{2}}$, and $\alpha=0$ GPa, as $\chi$ varies. Increasing values of the conservative component of $\mathbf{z}$ have a stabilizing effect. If $\chi=3$ GPa, instability occurs when $k_{0}=0.01$ $\mathrm{GPa/m^{2}}$, while stability is recovered for $k_{0}=0.1$ $\mathrm{GPa/m^{2}}$.
Table 5: Spectrum of $A$ for some values of $\chi$, with $k_{0}=0.01$, $\mathrm{GPa/m^{2}}$, $\varsigma=0.178\,\mathrm{GPa\cdot s/m^{2}}$, and $\alpha=0$ GPa.
$k_{0}=0.01$ $\mathrm{GPa/m^{2}}$, $\varsigma=0.178\,\mathrm{GPa\cdot s/m^{2}}$, $\alpha=0$ GPa.
$\chi$ $[\mathrm{GPa}]$ $0.5$ $1$ $1.7$ $1.9$
$\sigma(A)$ $[\mathrm{rad/s}]$ $-0.003+6492\,i$ $-0.01+6492\,i$ $-0.04+6492\,i$ $-0.05+6492\,i$
$-0.003-6492\,i$ $-0.01-6492\,i$ $-0.04-6492\,i$ $-0.05-6492\,i$
$-0.41$ $-0.39$ $-0.34$ $-0.32$
$-0.01+3570\,i$ $-0.04+3570\,i$ $-0.12+3570\,i$ $-0.16+3570\,i$
$-0.01-3570\,i$ $-0.04-3570\,i$ $-0.12-3570\,i$ $-0.16-3570\,i$
$-0.01+3570\,i$ $-0.04+3570\,i$ $-0.12+3570\,i$ $-0.16+3570\,i$
$-0.01-3570\,i$ $-0.04-3570\,i$ $-0.12-3570\,i$ $-0.16-3570\,i$
$-0.28$ $-0.22$ $-0.05$ +$0.01$
$-0.28$ $-0.22$ $-0.05$ +$0.01$
Table 6: Spectrum of $A$ for some values of $\chi$, with $k_{0}=0.1$, $\mathrm{GPa/m^{2}}$, $\varsigma=0.178\,\mathrm{GPa\cdot s/m^{2}}$, and $\alpha=0$ GPa.
$k_{0}=0.1$ $\mathrm{GPa/m^{2}}$, $\varsigma=0.178\,\mathrm{GPa\cdot s/m^{2}}$, $\alpha=0$ GPa.
$\chi$ $[\mathrm{GPa}]$ $0.5$ $1$ $1.9$ $3$
$\sigma(A)$ $[\mathrm{rad/s}]$ $-0.003+6493\,i$ $-0.013+6493\,i$ $-0.047+6493\,i$ $-0.117+6493\,i$
$-0.003-6493\,i$ $-0.013-6493\,i$ $-0.047-6493\,i$ $-0.117-6493\,i$
$-0.911+0\,i$ $-0.892+0\,i$ $-0.824+0\,i$ $-0.683+0\,i$
$-0.011+3570\,i$ $-0.043+3570\,i$ $-0.156+3570\,i$ $-0.388+3570\,i$
$-0.011-3570\,i$ $-0.043-3570\,i$ $-0.156-3570\,i$ $-0.388-3570\,i$
$-0.011+3570\,i$ $-0.043+3570\,i$ $-0.156+3570\,i$ $-0.388+3570\,i$
$-0.011-3570\,i$ $-0.043-3570\,i$ $-0.156-3570\,i$ $-0.388-3570\,i$
$-0.785+0\,i$ $-0.721+0\,i$ $-0.496+0\,i$ $-0.031+0\,i$
$-0.785+0\,i$ $-0.721+0\,i$ $-0.496+0\,i$ $-0.031+0\,i$
We fix $\chi=0.5$ GPa, $\alpha=0$ GPa, $\varsigma=0.178\,\mathrm{GPa\cdot s/m^{2}}$, and analyze the real psudospectrum of $A$ as $k_{0}$ varies. Results are in Figure 6: as $k_{0}$ grows, the three real eigenvalues of $A$ decrease; stability is assured.
Figure 6: Simultaneous presence of conservative and dissipative components of $\mathbf{z}$: real pseudospectrum of $A$ as a function of $k_{0}$.
If we allow $\varsigma$ to vary, by changing only $\mathfrak{N}$, and fixing $\chi=0.4$ GPa, $\alpha=0$ GPa, $k_{0}=0.09$ $\mathrm{GPa/m^{2}}$, as $\varsigma$ grows the real eigenvalues of $A$ tend toward 00 and, at the same time, the real part of the complex eigenvalues diminishes (see Figure 7). In this case, stability is assured for every value of $\mathfrak{N}$ considered.
Figure 7: Simultaneous presence of conservative and dissipative components of $\mathbf{z}$: real pseudospectrum of $A$ as $\mathfrak{N}$ varies.
Eventually, we allow $\chi$ to vary, keeping all the other parameters fixed. We set $\varsigma=0.178\,\mathrm{GPa\cdot s/m^{2}}$, $\alpha=0$ GPa, and $k_{0}=0.01$ $\mathrm{GPa/m^{2}}$. As $\chi$ grows, the complex eigenvalues of $A$ move towards the half-plane with positive real part and for $\chi>\sim 1.8$ GPa there is a Hopf-type transition towards instability. By increasing $k_{0}$, instability occurs for higher values of $\chi$; for example, for $k_{0}=0.1$ $\mathrm{GPa/m^{2}}$ instability occurs when $\chi>3$ GPa, see table 6.
Figure 8: Simultaneous presence of conservative and dissipative components of $\mathbf{z}$: real pseudospectrum of $A$ as $\chi$ varies.
8 Concluding remarks
When we analyze the dynamics of complex bodies, especially when microstructural diffusion is present, the evolution of waves with time-dependent amplitude usually leads to (and recurrently does so) non-normal matrices. The parameterized spectral analysis, which appears natural when some constitutive coefficients are experimentally unknown or are affected by uncertainties, may not be sufficient to clearly reveal stability and possible transition to unstable states. Exploring pseudospectra gives more reliable estimates, as those obtained in the present paper in the case of quasicrystals, for which constitutive parameters pertaining to the phason self-action are still unknown; their possible range can be determined by an experimental campaign aimed at exploring the possible stability/instability behavior indicated theoretically by the above results. The same question emerges when we homogenize at continuum scale manufactured lattice-like structures at microscopic or mesoscopic scale (like in the case of printed metamaterials): a single lattice leads generically to a classical continuum scheme; multiple lattices imply often multi-field continuum schemes, those falling within the general model-building framework of the mechanics of complex materials (an offspring of this framework is what has been discussed here). In the last case, the occurrence of genuinely non-normal matrices in the material stability analysis seems to be more common. In all these case, recourse to pseudospectra is helpful if not necessary and requires the following steps:
  • First, generic (complex) pseudospectra should be examined, for a range of values of $\varepsilon$. If the eigenvalues are found to be stable under such perturbations, they will be so also for real and for structured perturbations involving only matrices which preserve the particular structure present (such as the block structure, the symmetry of the blocks, etc.).
  • If, on the other hand, the (generic) pseudospectra indicates potential instability of the spectrum, one should compute and analyze the structured pseudospectra.
  • In particular, the evaluation of the spectral pseudo-abscissa, namely the maximum real part of the points in the pseudospectra, is especially important from a robust stability point of view (see, e.g., [15]). Given a stable matrix, the minimal (with respect to the norm) perturbation that implies a pseudospectrum trespassing into the unstable region characterizes the type of stability pertaining to the original matrix.
While in the special cases investigated in this paper we found that the pseudospectral analysis confirmed and validated the stability information obtained on the basis of the spectral analysis, in other cases this may not be necessarily so.
Acknowledgements. This research falls within the activities of the Theoretical Mechanics Research Group of the Mathematics Research Center E. De Giorgi of the Scuola Normale Superiore in Pisa. We acknowledge the support of GNFM-INDAM and GNCS-INDAM.